# GimbalDiffusion: Gravity-Aware Camera Control for Video Generation

### v1

@@ Abstract
Recent progress in text-to-video generation has achieved remarkable realism, yet fine-grained control over camera motion and orientation remains elusive. Existing approaches typically encode camera trajectories through relative or ambiguous representations, limiting explicit geometric control. We introduce GimbalDiffusion, a framework that enables camera control grounded in physical-world coordinates, using gravity as a global reference. Instead of describing motion relative to previous frames, our method defines camera trajectories in an absolute coordinate system, allowing precise and interpretable control over camera parameters without requiring an initial reference frame. We leverage panoramic $360^{\circ}$ videos to construct a wide variety of camera trajectories, well beyond the predominantly straight, forward-facing trajectories seen in conventional video data. To further enhance camera guidance, we introduce null-pitch conditioning, an annotation strategy that reduces the model’s reliance on text content when conflicting with camera specifications (e.g., generating grass while the camera points towards the sky). Finally, we establish a benchmark for camera-aware video generation by rebalancing SpatialVID-HQ for comprehensive evaluation under wide camera pitch variation. Together, these contributions advance the controllability and robustness of text-to-video models, enabling precise, gravity-aligned camera manipulation within generative frameworks.
“A serene aerial view captures a boat drifting down a dark river, framed by lush greenery and shimmering water under a peaceful sky.”
Refer to caption Refer to caption Refer to caption
“A sleek, modern home with high ceilings, natural light, and contemporary furnishings features a man in a blue suit in its open-plan living area.”
Refer to caption Refer to caption Refer to caption
Reference Ours AC3D [bahmani2025ac3d]
Figure 1: Reproducing real scenes using our representation. From reference frames from a driving video (left), GimbalDiffusion can be controlled to replicate its camera viewpoint with respect to gravity (center). In contrast, existing methods (right) do not accurately match the absolute reference pose.
@@ 1 Introduction
Text-to-video generation is rapidly advancing, offering astonishing photorealism and new creative workflows. Yet, the ability to control generated video content—especially camera movement and trajectory—remains limited. Early approaches rely solely on textual input which does not provide fine-grained control, such as specifying geometric parameters for camera paths and angles.
@@
Recent methods [he2024cameractrl, he2025cameractrl, bahmani2025vd3d, bahmani2025ac3d, bai2025recammaster, wan2025] address this limitation by conditioning video models with explicit camera trajectories, frequently encoded as Plücker rays. However, these representations are inherently ambiguous and lack the precision of real camera systems. Typically, camera extrinsics used for Plücker conditioning are defined relative to the first frame, making it impossible to explicitly control camera orientation with an absolute reference within the environment (e.g., with respect to gravity).
@@
In this work, we introduce GimbalDiffusion, a novel gravity-centric approach to data annotation that overcomes this challenge and enables precise, absolute user control over generated camera trajectories, as illustrated in LABEL:fig:Teaser. The name reflects the role of a gimbal—a device that stabilizes orientation in physical space—mirroring our use of gravity as a global reference to stabilize and disambiguate camera rotation. To resolve rotational ambiguity, we leverage geometrically calibrated 360° panoramic videos. By sampling camera crops over the sphere, we can remove the camera bias in human-captured datasets, allowing rare camera shots such as Dutch angles (45° roll) or extreme pitches to be sampled. This alignment enables the generation of cinematic shots, such as low- and high-angle perspectives, which are uncommon in regular video datasets. The effectiveness of this approach is demonstrated in fig. 1, where our method produces frames with camera views that more closely match the intended trajectories than previous methods.
To design our dataset from panoramic $360^{\circ}$ videos, original translation paths are preserved while camera rotations are randomly sampled. This lets our model learn a wide variety of camera trajectories, far beyond the predominantly straight, forward-facing paths seen in conventional video data.
@@
We further observe that text and camera angles are inherently entangled, and that care must be taken to ensure that appropriate control is achieved. For example, when asked to generate a video of a “green field under a blue sky” with a camera looking down, a model must learn to render grass, even if not specified in the prompt. To this end, we propose a null-pitch conditioning approach, where captioning is performed on forward-facing crops rather than the actual camera point of view.
@@
Finally, we propose a new evaluation benchmark for video generation with extreme camera angles. To do so, we use a large-scale dataset of high-quality internet videos, SpatialVID-HQ [wang2025spatialvidlargescalevideodataset], and rebalance the camera pitch to obtain diverse test samples. We will release this test dataset upon publication.
@@
  • A strategy named null-pitch conditioning for disentangling text from absolute camera angle;
  • A novel benchmark for video generation with extreme camera angles.
Figure 2: Training data pipeline. (a) We extract the camera poses from a dataset of 360° panoramic videos in equirectangular format (see Sec. 3.2). (b) We randomly sample a camera rotation trajectory on the sphere, from which we generate field of view masks in equirectangular format for both the projected image and the null-pitch conditioning (see Sec. 3.3). (c) We integrate the two previous steps by combining the sampled rotation with the estimated camera poses. This process creates a dataset of perspective images with their camera poses, along with null-pitch reference images. The main advantage of this approach is the ability to sample camera viewing directions across the entire sphere, in contrast to using natural human-captured videos, which are heavily biased towards null roll and pitch. Additionally, we generate a text description for both the perspective video and the null-pitch video using a VLM.
@@ 2 Related Work
Generative video modeling has evolved rapidly, from early GAN-based methods that produced short, low-resolution clips [tulyakov2018mocogan, li2018video, Acharya2018] to large-scale diffusion-based architectures capable of generating temporally coherent, semantically consistent videos [Weng2024]. Recent open-source models such as CogVideoX [yang2024cogvideox] and WAN [wan2025] push the frontier of video length, resolution, and controllability while proposing both text-to-video and image-to-video variants.
Editing generated videos remains challenging: precise modification to content, viewpoint, or lighting is difficult once frames are synthesized. To address this, several control strategies have emerged, introducing user-guided conditioning based on edge maps, depth, sketches, or explicit 3D geometry for structural control [chen2023controlavideo, Yenphraphai2024ImageSculpting, chang2024magicpose, hu2023animateanyone], and illumination-aware guidance for lighting control [kocsis2024lightit, Magar_lightlab]. These mechanisms allow precise control over motion, composition, and scene appearance, substantially improving the usability and compositional fidelity of generative video systems.
@@
Beyond content and appearance, viewpoint control has become key in generative video. A common representation for such control is Plücker rays, described in Sec. 3.1. This formulation enables differentiable, geometrically consistent camera manipulation.
In static image generation, PreciseCam [bernal2025precisecam] introduces absolute control over camera pose by conditioning on roll, pitch, field of view, and lens distortion alongside textual prompts. For image-to-video generation, methods such as MotionCanvas [xing2025motioncanvas], Diffusion-as-Shader [gu2025das], and TrajectoryCrafter [mark2025trajectorycrafter] employ point cloud conditioning to maintain spatial coherence while navigating a scene from an initial reference image. While effective, these approaches depend on the depth and pose estimation quality of the initial reference frame, which can restrict flexibility and generalization.
In text-to-video synthesis, AC3D [bahmani2025ac3d] allows user-defined camera trajectories but conditions motion relative to the first frame only, leaving global rotational and translational ambiguities unresolved. Similarly, Li et al. [li2025cameras] manipulate positional encodings to control viewpoint changes, yet still lack a consistent absolute reference such as gravity.
These works underscore the need for representations that enable precise, absolute camera control and scalable trajectory generation across space and time. Our approach addresses this gap by aligning camera orientation with gravity and anchoring motion within a consistent world frame, advancing controllable text-to-video synthesis.
@@
Estimating accurate camera poses from images is a long-standing problem in 3D vision. Classical structure-from-motion pipelines such as VisualSFM [wu2011visualsfm], OpenMVG [moulon2016openmvg], and COLMAP [Schonberger2016COLMAP] reconstruct camera intrinsics and extrinsics through feature matching, triangulation, and global bundle adjustment. More recent learning-based approaches jointly infer scene geometry and pose from data: VGGT [wang2025vggt] leverages transformer-based global geometry tokens for pose regression, while DUSt3R [wang2024dust3r] and MASt3R [leroy2024grounding] predict dense correspondence fields to recover metrically scaled 3D structure. MegaSAM [Jiang2024MegaSAM] extends these ideas to foundation scale, achieving state-of-the-art alignment accuracy.
Despite their accuracy, these systems typically operate in arbitrary coordinate frames and lack physical grounding in gravity or a global world reference, limiting their absolute interpretability. To mitigate this, several works estimate the gravity direction from single images [jin2022PerspectiveFields, veicht2024geocalib, xian2019uprightnet]. However, acquiring datasets with gravity annotations often requires the use of auxiliary sensors, such as accelerometers, or additional semantic information about the scene. Examples include EDINA [Do_2022_EgoSceneMSR], Horizon Lines in the Wild [workman2016hlw], and the KITTI Horizon dataset [kluger2020temporally].
To circumvent such constraints, we leverage geometrically calibrated 360° panoramic videos from WEB360 [huang2025vipe] as backbone for our data generation. By combining the up vector derived from the panoramic video with SfM-estimated intrinsics and extrinsics, we obtain a consistent absolute reference for gravity-aware camera control.
Figure 3: Training data samples from our data augmentation pipeline, capturing a highly diverse set of rotation trajectories from 360° videos. Both the trajectory and the prompt are generated on-the-fly in our dataloader. Each color in the graph corresponds to a video above.
@@ 3 Method
We aim to train a camera-conditioned model on precise, gravity-referenced camera viewpoints. We refer to this global gravity reference as an absolute reference, in contrast to existing relative methods that describe camera motion with respect to a previous frame. Because our absolute reference is aligned with gravity, pitch and roll are well defined, whereas yaw remains unconstrained. Throughout the paper, absolute rotation therefore denotes pitch and roll expressed in this shared gravity-aligned reference.
However, no large-scale video dataset provides the gravity annotations we need to train such a model. This section outlines our pipeline for obtaining such data, summarized in fig. 2. In fig. 2 (a) we extract camera pose annotations from a dataset of 360° panoramic videos, as detailed in Sec. 3.2. Next, in (b), we sample a camera trajectory as described in Sec. 3.2. Finally, in (c), we combine the two previous steps to produce our dataset. We repeat the same procedure for an additional version of each trajectory with null pitch and roll, described in Sec. 3.3. Before detailing the full pipeline, the following section explains the standard representation to encode camera poses in neural-network-compatible buffers.
@@ 3.1 Representing cameras
Given a camera extrinsic matrix $\mathbf{E}_{f}\!=\![\mathbf{R}_{f}|\mathbf{t}_{f}]$ and intrinsic matrix $\mathbf{K}_{f}$, a common strategy for conditioning video generation models on viewpoint is to use Plücker rays [kant2024spad, he2024cameractrl, he2025cameractrl, bahmani2025vd3d, bahmani2025ac3d]. For a conditioning frame $f$, we can define a per-pixel map of Plücker rays as $\mathbf{p}_{f,u,v}\!=\!(\mathbf{t}_{f}\times\mathbf{d}_{f,u,v}^{\prime},\mathbf{d}_{f,u,v}^{\prime})$, where $\mathbf{t}_{f}$ denotes the camera origin in world coordinates, and $\mathbf{d}_{f,u,v}$ is the direction vector from the camera center toward the center of pixel $(u,v)$. Each pixel is thus represented as a 6D vector describing a ray in 3D space. The ray direction can be computed as $\mathbf{d}_{f,u,v}\!=\!\mathbf{R}_{f}\mathbf{K}_{f}^{-1}[u,v,1]^{T}+\mathbf{t}_{f}$, with $\mathbf{d}_{f,u,v}^{\prime}$ denoting the normalized direction vector $\mathbf{d}_{f,u,v}$.
@@ 3.2 Creating varied camera paths from 360° videos
Most video datasets exhibit a significant bias towards forward-looking views, where the horizon remains level or close to the center of the image, as shown in fig. 4. However, such “normal” viewpoints restrict the expressive range that a filmmaker might wish to achieve in a storyboard. For instance, these datasets often lack cinematic shots, such as top-down, low-angle, or Dutch-angle shots. As a result, generative video models trained on such data struggle to produce these types of cinematic shots, since they fall outside the training distribution.
To address this limitation, we propose generating synthetic camera trajectories by sampling from 360° videos and rectifying them so that the horizon position is known throughout each sequence.
@@
From a 360° video, we compute the relative motion between each frame by extracting six perspective crops (front, back, left, right, top, bottom), and processing the resulting video with structure-from-motion (SfM). This results in a set of camera poses $\mathbf{E}_{\text{SfM},f}$ at each frame $f$. Taking the first frame of the sequence $f=0$ as reference, we can then compute the relative motion
@@
We can generate several perspective videos from a 360° sequence by sampling random camera rotation paths. We outline our approach in Algorithm 1. We first sample random pitch, roll, and yaw angles uniformly to obtain an initial viewpoint. Next, 1 to 4 additional random rotations are sampled and distributed across the sequence length. Those rotations are applied sequentially to produce smooth and varied motion paths.

### v2

@@ Abstract
Recent progress in text-to-video generation has achieved remarkable realism, yet fine-grained control over camera motion and orientation remains elusive, especially with extreme trajectories (e.g., a 180° turnaround, or looking directly up or down). Existing approaches typically encode camera trajectories using relative or ambiguous representations, limiting precise geometric control and offering limited support for large rotations. We introduce GimbalDiffusion, a framework that enables camera control grounded in physical-world coordinates, using gravity as a global reference. Instead of describing motion relative to previous frames, our method defines camera trajectories in an absolute coordinate system, allowing accurate, interpretable control over camera parameters. Using panoramic $360^{\circ}$ videos for training, we cover the full sphere of possible viewpoints, including combinations of extreme pitch and roll that are out-of-distribution of conventional video data. To improve camera guidance, we introduce null-pitch conditioning, a strategy that prevents the model from overriding camera specifications in the presence of conflicting prompt content (e.g., generating grass while the camera points toward the sky). Finally, we propose new benchmarks to evaluate gravity-aware camera-controlled video generation, assessing models’ ability to generate extreme camera angles and quantify their input prompt entanglement.
@@ 1 Introduction
Text-to-video generation is rapidly advancing, offering increasingly photorealistic results and new creative workflows. Yet, the ability to control generated video content—especially camera trajectory—remains limited, particularly for extreme camera rotations such as top-down shots, Dutch angles, or full 360° pans. Early approaches rely solely on textual input which does not provide fine-grained control, such as specifying geometric parameters for camera paths and angles.
@@
Recent methods [13, 14, 3, 2, 4, 34] address this limitation by conditioning video models with explicit camera trajectories, frequently encoded as Plücker rays. However, these representations are inherently ambiguous and lack the precision of real camera systems. Typically, camera extrinsics used for Plücker conditioning are defined relative to the first frame, making it impossible to explicitly control camera orientation with an absolute reference within the environment (e.g., with respect to gravity). As a result, these methods are limited to modest rotations and fail at extreme camera angles.
@@
In this work, we introduce GimbalDiffusion, a novel gravity-centric approach to data annotation that overcomes this challenge and enables precise, absolute user control over generated camera trajectories, as illustrated in LABEL:fig:Teaser. The name reflects the role of a gimbal—a device that stabilizes orientation in physical space—mirroring our use of gravity as a global reference to stabilize and disambiguate camera rotation. To resolve rotational ambiguity, we leverage geometrically calibrated 360° panoramic videos. By sampling camera crops over the sphere, we can remove camera bias in human-captured datasets, enabling the sampling of rare cinematic camera shots, such as barrel rolls (a “twisting” camera rotation), extreme low- and high-angle perspectives, or looping 360° rotations. The effectiveness of this approach is demonstrated in fig. 1, where our method produces frames whose camera views more closely match the intended trajectories than those of previous methods. This enables our model to learn a wide range of camera trajectories, far beyond the predominantly straight, forward-facing paths in conventional videos.
Forward Up Back Down Forward
Ours
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
UCPE
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Gen3C
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A spacious living room with a gray sofa facing a wooden coffee table. A large window behind sheer curtains lets soft afternoon light enter the room. A bookshelf filled with books lines one wall, and a floor lamp stands near the sofa.”
Figure 1: Extreme rotation handling. Here, the camera starts forward and performs a motion akin to a “backflip”: looking straight up, then backwards, down, and back to the starting position. Our method (top) precisely follows the desired camera trajectory while accurately reproducing the original content from the first frame. Other methods, like UCPE [50] and GEN3C [30] (initialized with PreciseCam [5]), fail at this task. Refer to the supp. for the full videos.
@@
We further observe that text and camera angles are inherently entangled, and that care must be taken to ensure that appropriate control is achieved. For example, when asked to generate a video of a “green field under a blue sky” with a camera looking down, a model must learn to render only grass, even if other concepts are specified in the prompt. To this end, we propose a null-pitch conditioning approach in which captioning is performed on forward-facing crops rather than on the actual camera point of view.
@@
Finally, we propose two new evaluation benchmarks. The first evaluates video generation with extreme camera angles by leveraging the large-scale, high-quality dataset SpatialVID-HQ [35] and rebalancing the camera pitch distribution to obtain diverse test samples. The second evaluates entanglement between prompt and camera angle by annotated crops of panoramas from the PolyHaven dataset [12]. We will release these test datasets upon publication.
@@
  • A strategy named null-pitch conditioning for disentangling text from absolute camera angle, critical for extreme viewpoints where prompt content and camera orientation diverge most;
  • Two evaluation benchmarks for: video generation with extreme camera angles, and assessing the entanglement between input prompt and camera orientation.
@@ 2 Related Work
Generative video modeling has evolved rapidly, from early GAN-based methods that produced short, low-resolution clips [32, 26, 1] to large-scale diffusion-based architectures capable of generating temporally coherent, semantically consistent videos [40]. Recent open-source models such as CogVideoX [46] and WAN [34] push the frontier of video length, resolution, and controllability while proposing both text-to-video and image-to-video variants.
Editing videos remains challenging: precise changes to content, viewpoint, or lighting are difficult once the frames have been synthesized. To address this, control strategies have emerged, introducing user-guided conditioning based on edge maps, depth, sketches, or explicit 3D geometry for structural control [7, 47, 6, 16], and illumination-aware guidance for lighting control [23, 28]. These mechanisms allow precise control over motion, composition, and scene appearance, enhancing the usability and compositional fidelity of generative video systems.
@@
In static image generation, PreciseCam [5] (similarly, [27]) provide text-to-image models with absolute control over camera pose by conditioning on roll, pitch, field of view, and lens distortion alongside textual prompts. For image-to-video generation, ViewCrafter [49] and GEN3C [30] generate novel views from a single input image, but cannot handle scene motion. Methods such as MotionCanvas [45], Diffusion-as-Shader [11], and TrajectoryCrafter [48] maintain spatial coherence with point cloud conditioning from an initial reference image. While effective, these approaches depend on the reference frame’s depth and pose estimation quality, limiting flexibility and generalization.
In text-to-video synthesis, AC3D [2] allows user-defined camera trajectories but conditions motion relative to the first frame only, leaving global rotational and translational ambiguities unresolved. Similarly, Li et al. [25] manipulate positional encodings to control viewpoint changes, yet still lack a consistent absolute reference such as gravity. Concurrent with our work, UCPE [50] extends positional encodings to encode a latitude up map. However, its data generation is complex and requires matching 360° videos with rotations extracted from a secondary dataset. It also cannot handle extreme camera angles (see fig. 1).
These works underscore the need for precise, absolute camera control that remains robust under extreme rotations. Our approach addresses this gap by aligning camera orientation with gravity and anchoring motion within a consistent world frame, allowing extreme camera rotations in text-to-video synthesis.
Panorama video generation methods
[37, 43, 51] adapt video generation backbones for the task of generating 360° equirectangular videos. In addition to text, they can be conditioned on an input image [37, 51], optical flow [37] and derotated flow [51]. While gravity-aligned camera control can be achieved by cropping the desired field of view out of the generated panorama, this comes with several disadvantages: 1) most of the panorama pixels are discarded, resulting in severe resolution degradation; 2) the desired concepts described in the prompt can be cropped out and thus not appear in the video; and 3) current text-conditioned methods [43, 37] do not provide control over camera translation. We demonstrate these downsides in the supplementary material.
@@
Estimating accurate camera poses from images is a long-standing problem in 3D vision. Classical structure-from-motion pipelines such as VisualSFM [42], OpenMVG [29], and COLMAP [31] reconstruct camera intrinsics and extrinsics through feature matching, triangulation, and global bundle adjustment. More recent learning-based approaches jointly infer scene geometry and pose from data: VGGT [36] leverages transformer-based global geometry tokens for pose regression, while DUSt3R [39] and MASt3R [24] predict dense correspondence fields to recover metrically scaled 3D structure. MegaSAM [18] extends these ideas to foundation scale, achieving state-of-the-art alignment accuracy. ViPE [17] extends these ideas to handle in-the-wild videos.
Despite their accuracy, these systems typically operate in arbitrary coordinate frames and lack physical grounding in gravity or a global world reference, limiting their absolute interpretability. To mitigate this, several works estimate the gravity direction from single images [19, 33, 44]. However, acquiring datasets with gravity annotations often requires auxiliary sensors, such as accelerometers, or additional semantic information about the scene. Examples include EDINA [9], Horizon Lines in the Wild [41], and the KITTI Horizon dataset [22]. ScanNet [8] proposes a dataset of 3D scenes captured using an RGB-D sensor, but exhibits limited diversity, primarily focusing on small indoor environments such as offices, apartments, and bathrooms.
To circumvent such constraints, we geometrically calibrate 360° videos from the PanoVid dataset [43]. By combining the up vector derived from the panoramic video with SfM-estimated intrinsics and extrinsics, we obtain a consistent absolute reference for conditioning video models on gravity-aware camera control.
Figure 2: Training data pipeline. On the left, we start by estimating the camera poses and gravity up vectors from a dataset of 360° panoramic videos in equirectangular format (see Sec. 3.2). We then randomly sample a camera rotation trajectory and camera intrinsics on the sphere, which are used for projecting the panoramic frames to perspective images. We integrate the two previous steps by combining the sampled rotation with the estimated camera poses (containing also translations). This process creates a dataset of perspective images (in green) with their absolute camera poses. The main advantage of this approach is the ability to sample camera viewing directions across the entire sphere, in contrast to using natural human-captured videos, which are heavily biased towards null roll and pitch, and minimal rotation amplitude. We also project the panorama on the rotation trajectory with pitch and roll removed (in blue), yielding the null-pitch video from which we obtain a textual description (see Sec. 3.3).
@@ 3 Method
We aim to train a camera-conditioned model that supports extreme camera rotations, including full 360° trajectories, extreme pitch, and large roll, while maintaining precise control. To achieve this, we ground camera poses in an absolute, gravity-aligned coordinate system. We refer to this global gravity reference as an absolute reference, in contrast to existing relative methods that describe camera motion with respect to a previous frame. Because our absolute reference is aligned with gravity, pitch and roll are well defined, whereas yaw remains unconstrained. Throughout the paper, absolute rotation therefore denotes pitch and roll expressed in this shared gravity-aligned reference.
However, no large-scale video dataset provides the gravity annotations we need to train such a model. This section outlines our pipeline for obtaining such data, summarized in fig. 2. We first extract camera pose annotations from a dataset of 360° panoramic videos (Sec. 3.2 and fig. 2, left). Next, we sample a camera trajectory (Sec. 3.2 and fig. 2, center-bottom). Finally, we combine the two previous steps to produce our dataset. We repeat the same procedure for an additional version of each trajectory with null pitch and roll (Sec. 3.3, fig. 2, center-top). Before detailing the full pipeline, Sec. 3.1 explains the standard representation to encode camera poses in neural-network-compatible buffers.
@@ 3.1 Representing cameras
Given a camera extrinsic matrix $\mathbf{E}_{f}\!=\![\mathbf{R}_{f}|\mathbf{t}_{f}]$ and intrinsic matrix $\mathbf{K}_{f}$, a common strategy for conditioning video generation models on viewpoint is to use Plücker rays [20, 13, 14, 3, 2]. For a conditioning frame $f$, we can define a per-pixel map of Plücker rays as $\mathbf{p}_{f,u,v}\!=\!(\mathbf{t}_{f}\times\mathbf{d}_{f,u,v}^{\prime},\mathbf{d}_{f,u,v}^{\prime})$, where $\mathbf{t}_{f}$ denotes the camera origin in world coordinates, and $\mathbf{d}_{f,u,v}$ is the direction vector from the camera center toward the center of pixel $(u,v)$. Each pixel is thus represented as a 6D vector describing a ray in 3D space. The ray direction can be computed as $\mathbf{d}_{f,u,v}\!=\!\mathbf{R}_{f}\mathbf{K}_{f}^{-1}[u,v,1]^{T}+\mathbf{t}_{f}$, with $\mathbf{d}_{f,u,v}^{\prime}$ denoting the normalized direction vector $\mathbf{d}_{f,u,v}$.
@@ 3.2 Creating varied camera paths from 360° videos
Most video datasets exhibit a significant bias toward forward-looking views with horizon level or near the center of the image, as shown in the supp. material. However, such “normal” viewpoints limit storytelling expressivity by lacking cinematic shots, such as top-down, low-angle, or barrel-roll shots. Consequently, generative video models trained on such data struggle to produce these types of cinematic shots, as they lie outside the training distribution (see fig. 1). To address this limitation, we propose generating synthetic camera trajectories by sampling and rectifying 360° videos.
@@
From a 360° video, we compute the relative motion between frames by extracting six perspective crops (front, back, left, right, top, bottom), and processing the resulting video with a pose estimation method [17]. This yields a set of camera poses $\mathbf{E}_{\text{SfM},f}$ for each frame $f$. Taking the first frame of the sequence $f=0$ as reference, we can then compute the relative motion
@@
Gravity annotation.
Pose estimation methods such as [17] provide relative estimates of camera rotations. To anchor the videos within a shared global reference frame, we align the relative camera rotations with gravity. Specifically, we use [19] to estimate the camera roll and pitch angles at frame 0. Because it operates on perspective images, we first project the panoramic frame into eight perspective views evenly sampled about the yaw axis at fixed 90° field of view. We then estimate the camera up vectors from these views and rotate them back to the null yaw reference frame. We then average the up vectors across these views, yielding $\mathbf{\tilde{u}}_{0}$. From this, we derive the absolute, gravity-aligned camera pose for each frame as
$\mathbf{t}_{\text{abs},f}=\psi(\mathbf{\tilde{u}}_{0})\mathbf{t}_{\text{rel},f}$
$\mathbf{R}_{\text{abs},f}=\mathbf{R}_{\text{pano},f}\psi(\mathbf{\tilde{u}}_{0})\mathbf{R}_{\text{rel},f}$ (2)
$\mathbf{E}_{\text{abs},f}=\varphi(\mathbf{R}_{\text{pano},0})[\mathbf{R}_{\text{abs},f}|\mathbf{t}_{\text{abs},f}]\;,$
where $\varphi(\cdot)$ is a function that preserves only the roll and pitch angles (removes the yaw) of the first sampled camera pose, and $\psi(\cdot)$ aligns camera poses to the gravity up vector. After this adjustment, the transformed pose $\mathbf{E}_{\text{abs},f}$ has no rotation about the global up-vector axis. Please refer to the supplementary material for the full definition of $\varphi(\cdot)$ and $\psi(\cdot)$. This process establishes a gravity-aligned coordinate system consistent across all videos, enabling rotations to be defined within a common global frame. Consequently, each sequence is initialized with null yaw, reducing ambiguity and facilitating training with this representation.
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“[…] an aerial view of a large resort complex featuring multiple buildings with red roofs […]”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“[…] large outdoor gathering of people and cars. […] The sky is overcast […]”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A young man in a black hoodie and pants walks along a […] street, engrossed in his smartphone. […]”
Videos and captions Camera poses
Figure 3: Training data samples from our data augmentation pipeline, capturing a highly diverse set of rotation trajectories from 360° videos. Both the trajectory and the prompt are generated on-the-fly in our dataloader. See the supp. for full videos.
@@
We can generate several perspective videos from a 360° sequence by sampling random camera rotation paths. We outline our approach in Algorithm 1. We first sample random pitch, roll, and yaw angles uniformly to obtain an initial viewpoint. Next, 1 to 3 additional random rotations are sampled and distributed across the sequence length. Those rotations are applied sequentially to produce smooth and varied motion paths.